# LFTC-25-11 / 105

### v1

@@ I Introduction
The phenomenological success of this particular approach in describing hadron structure is noteworthy. Its results are comparable to those obtained using a variety of other well-established methods, including, Schwinger-Dyson methods Roberts and Williams (1994); Roberts and Schmidt (2000), QCD sum rules Braguta and Onishchenko (2004); Aliev and Savci (2004); Aliev et al. (2009), AdS/QCD frameworks Brodsky and de Teramond (2008), Effective Field Theory (EFT) Bijnens et al. (1998), Covariant light-front dynamics Karmanov (1996); Leão and de Melo (2023), and also Point-form quantum mechanics Biernat and Schweiger (2014).
@@
However, LFQFT is not without challenges. A critical issue is the loss of covariance in certain physical processes de Melo and Frederico (1997); de Melo et al. (1999); Bakker et al. (2002); Choi and Ji (2004). To ensure the full covariance of the electromagnetic current, it is necessary to go beyond the simple valence component and incorporate non-valence contributions (often called zero-modes) into the electromagnetic current’s matrix elements de Melo and Frederico (1997); de Melo et al. (1999); de Melo and Frederico (2012); De Melo (2019); Li et al. (2019).
@@
In a recent publication Zhang et al. (2025), specifically concerning the Yukawa model for the nucleon, the problematic inclusion of zero modes ($k^{+}=0$ modes, and non-valence terms, higher Fock-sector contributions, often referred to as pair terms) was discussed. The dependence of the obtained results on the adopted reference frames was explored, highlighting the necessity of these extra contributions to recover the expected Lorentz invariance and frame-independence of physical observables, such as form factors.
In the LF approach, the matrix elements of the current operator $\hat{J}^{\mu}$ are typically calculated using the overlap of Light-Front Wave Functions (LFWFs). For the full covariance to hold, the results derived from different current components, $\langle P^{\prime}|\hat{J}^{+}|P\rangle$ and $\langle P^{\prime}|\hat{J}^{-}|P\rangle$, must be consistent. The necessity to include zero modes or non-valence contributions, is often a consequence of the truncation of the Fock space or the specific component of the current employed. The authors in Zhang et al. (2025) (and references therein) investigate how the proper treatment of these terms is essential for the cancellation of spurious frame dependence, a clear signature of the restoration of Poincaré symmetry de Melo et al. (1998, 1999).
@@
Light-front dynamics is inherently a constrained system, and thus precludes direct canonical quantization. This non-canonical structure introduces complications, particularly in fermionic theories, where certain components of the spinor fields are not independent degrees of freedom; they are subject to constraint equations, and these are often referred to as bad components. These bad components necessarily enter into the definition of the electromagnetic operators, subsequently leading to modified, potentially non-standard currents. Despite the fact that the simplest Ward-Takahashi identity connects the electromagnetic vertex and the propagator, the primary unanswered question involves the integral relationship tying the current to the vertex.
@@
The current investigation significantly advances the initial results by exploring a more generalized scope (or a broader class of systems/parameters). Central to this expanded analysis is a detailed examination of the Ward-Takahashi (WT) identities as they apply to the fermionic sector of the theory. Crucially, the analysis reaffirms the persistent requirement to incorporate non-valence contributions—those originating outside the simplest quasi-particle picture, to accurately model the electromagnetic current.
@@
The present article is organized as follows. In Sec.II, we provide a proof for the Ward-Takahashi identity at one loop in the covariant formulation for the given system. It is followed by the proof in the LF formulation in Sec.III. In Sec.III.1, the LF expression for self-energy is obtained by integrating out the energy component in the loop momentum. The contribution of zero modes to this expression is explained in the process. Further, the LF expression for the electromagnetic current is similarly obtained in Sec.III.2 and the LF Ward-Takahashi identity is established. Sec.IV contains a few concluding remarks. The conventions used throughout the article, and a few useful identities are provided in Appendix A, and a part of the algebra from Sec.III.2 is relegated to Appendix B.
@@ II Ward-Takahashi identity in the covariant formulation
Consider a spin-0 bound-state made up of a fermion-antifermion pair, for example a pion, given by a $\gamma^{5}$-type of an interaction vertex. The self-energy diagram for such a state is shown in Fig.(1).
@@
The expression for the self-energy of this bound-state is given in the equal-time covariant formulation by
@@
Now consider the electromagnetic current corresponding to this bound state. It is represented by the Feynman triangle diagram shown in Fig.(2).
@@ III Ward-Takahashi identity in the LF formulation
The light-front formulation is usually implemented in the form a Hamiltonian field theory. Thus, all particles in the theory are on-shell. We now establish the Ward-Takahashi identity at one-loop for the system in Sec.II i.e. a fermion-antifermion bound-state of spin 0, but in the LF formulation. For this, we consider the 4-D loop integral in the covariant theory and integrate out the energy component from the loop integral. This procedure is essentially equivalent to calculating the time-ordered graphs of the theory in old-fashioned Hamiltonian perturbation theory (Bhamre and Misra, 2022). While establishing the Ward-Takahashi identity such, one needs to be mindful of the zero modes that arise when the longitudinal LF momentum vanishes. We proceed to show this below.
@@ III.1 The self-energy
where $k^{-}_{\text{on}}=\frac{k_{\perp}^{2}+m^{2}}{2k^{+}}$ is the on-shell LF energy of a particle with momentum $k^{\mu}=(k^{+},k^{-},k^{\perp})$. Thus, a fermion propagator can be split into an ‘on-shell’ term and an ‘off-shell’ term as follows
@@
the on-shell $4$-momentum being given by $k^{\mu}_{\text{on}}=(k^{+},k^{-}_{\text{on}},k^{\perp})$. The ‘off-shell’ term is a non-propagating term which is typically present in the LF formulation. Similarly, the other propagator in the fermion loop is written as
@@
Using the properties of gamma matrices $\{\gamma^{5},\gamma^{\mu}\}=0$ and $(\gamma^{+})^{2}=0$, it is seen that the last last term of the above integrand vanishes. Also, the traces in the remaining terms can be calculated using the properties of gamma matrices, giving
@@
$I_{1}^{-}$ in Eq.(16) has poles at $k_{1}^{-}=\frac{f_{1}-i\epsilon}{2k^{+}}$ and $k_{2}^{-}=P^{-}-\frac{f_{2}-i\epsilon}{2(P^{+}-k^{+})}$. Note that there is no $k^{-}$-dependence in the numerator as the $4$-momenta appearing there ($k^{\mu}_{\text{on}}$, $(P-k)^{\mu}_{\text{on}}$) are on-shell momenta. For $k^{+}<0$, both the poles $k_{1}^{-}$ and $k_{2}^{-}$ lie above the real axis while for $k^{+}>P^{+}$ , both lie below it. Hence, closing the contour on the opposite side of the position of the pole, the integral vanishes for each of these two ranges of the longitudinal momentum. For $0<k^{+}<P^{+}$, $k_{1}^{-}$ lies below the real axis whereas $k_{2}^{-}$ lies above. Closing the contour below, the residue theorem gives
@@
However, this is not the case in Eqs.(17,18). $I_{2}^{-}$ in Eq.(17) has a pole at $k_{1}^{-}=\frac{f_{1}-i\epsilon}{2k^{+}}$ which does go to infinity as $k^{+}\to 0$. This is a purely divergent zero-mode term which manifests here as an end-point singularity. In order to correctly account for this pole at infinity, we make a change of variable so as to bring the pole to the origin, regularize it, and eventually take the regulator to zero (Bhamre and Misra, 2022). Let $k^{-}=\frac{1}{u}$ in Eq.(17). Then,
@@
The first integral above has poles at $u_{1}=-i\delta$ and $u_{2}=\frac{2k^{+}}{f_{1}-i\epsilon}$. Thus, the integral is non-vanishing only for $k^{+}>0$. Similarly, the second integral contributes only through the region $k^{+}<0$. Applying the residue theorem, and then taking the regulator $\delta$ to zero, we get
@@
Considering $P^{\prime+}>P^{+}$, the integration regions of the log-divergent $k^{+}$-integrals combine to give
@@ III.2 The electromagnetic current
The $P$-dependent and $P^{\prime}$-dependent parts of $q_{\mu}J^{\mu}(P,P^{\prime})$ in Eq.(37) separate out and are contained in the two sets of integrals $I^{+}_{1}$ through $I^{+}_{6}$, and $I^{+}_{7}$ through $I^{+}_{12}$ respectively. Substituting the sum over all $I^{+}_{i}$ in Eq.(37) eventually gives (algebra worked out in Appendix B)
@@
Leaving the zero-mode contribution in Eq.(38) aside, it is worth noting that there are two regions of $k^{+}$-integration - one from 0 to $P^{+}$, and another from $P^{+}$ to $P^{\prime+}$. Considered as a time-ordered graph in LF, the diagram in Fig.(2) can only contribute to the region $0<k^{+}<P^{+}$ due to 3-momentum conservation and the fact that the longitudinal momentum component in LF can only be non-negative. The contribution to the rest of the longitudinal momentum region i.e. $P^{+}<k^{+}<P^{\prime+}$ is provided by the other time-ordered LF diagram, viz. the ‘pair production’ diagram shown in Fig.(3). Thus, without the pair-production diagram, the Ward-Takahashi identity in LF cannot be established.
@@ IV Conclusion
The preceding analysis delivers a rigorous, one-loop proof of the Ward-Takahashi identity for a spinless, fermionic bound-state, comparing and contrasting its realization within both the covariant and Light-Front formulations of quantum field theory.
@@
By correctly incorporating these end-point singularities, the contributions typically associated with zero modes - the dynamic complexity unique to the light-front - are naturally and elegantly taken into account in the proof. This demonstrates that the WT identity, and thus gauge invariance, is preserved in the LF theory without the need for ad hoc zero-mode operators, provided the loop integrals are handled with scrupulous care. Along with the ‘regular’ diagram viz. the triangle diagram for the electromagnetic current, the ‘pair production’ diagram is also necessary to establish the proof of the identity. This is consistent with the fact that the proof is achieved in the Hamiltonian time-ordered perturbation theory on the light-front.
This work not only validates the Light-Front Hamiltonian time-ordered perturbation theory as a consistent framework for handling gauge theories but also provides a clear, instructive procedure for translating between covariant and LF results. The explicit connection made between end-point singularities in loop integrals and the required physical contributions (like pair production diagrams) significantly enhances our understanding of how fundamental symmetries, like gauge invariance, are preserved in non-trivial field-theoretic contexts within the light-front framework. This rigorous verification strengthens the theoretical foundation for future applications of Light-Front Quantization to more realistic and complex bound-state problems in particle and nuclear physics.
Acknowledgements
DB acknowledges the support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil, Process No. 152348/2024-7. JPBCM acknowledges the support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), Process No. 2023/09539-1, Instituto Nacional de Ciência e Tecnologia-Nuclear Physics and Applications (INCT-FNA, MCTI), Brazil, Process No.464898/2014-5, and FAPESP, Brazil, thematic project, No. 2024/17816-8.

### v2

@@ I Introduction
The phenomenological success of this particular approach in describing hadron structure is noteworthy. Its results are comparable to those obtained using a variety of other well-established methods, including, Schwinger-Dyson methods Roberts and Williams (1994); Roberts and Schmidt (2000), QCD sum rules Braguta and Onishchenko (2004); Aliev and Savci (2004); Aliev et al. (2009), anti-de Sitter (AdS)/QCD frameworks Brodsky and de Teramond (2008), effective field theory (EFT) Bijnens et al. (1998), covariant light-front dynamics Karmanov (1996); Leão and de Melo (2023), and also point-form quantum mechanics Biernat and Schweiger (2014).
@@
However, LFQFT is not without challenges. A critical issue is the loss of covariance in certain physical processes de Melo and Frederico (1997); de Melo et al. (1999); Bakker et al. (2002); Choi and Ji (2004). To ensure the full covariance of the electromagnetic current, it is necessary to go beyond the simple valence component and incorporate nonvalence contributions (often called zero modes) into the electromagnetic current’s matrix elements de Melo and Frederico (1997); de Melo et al. (1999); de Melo and Frederico (2012); De Melo (2019); Li et al. (2019).
@@
In a recent publication Zhang et al. (2025), specifically concerning the Yukawa model for the nucleon, the problematic inclusion of zero modes ($k^{+}=0$ modes, and nonvalence terms, higher Fock-sector contributions, often referred to as pair terms) was discussed. The dependence of the obtained results on the adopted reference frames was explored, highlighting the necessity of these extra contributions to recover the expected Lorentz invariance and frame independence of physical observables, such as form factors.
In the LF approach, the matrix elements of the current operator $\hat{J}^{\mu}$ are typically calculated using the overlap of light-front wave functions. For the full covariance to hold, the results derived from different current components, $\langle P^{\prime}|\hat{J}^{+}|P\rangle$ and $\langle P^{\prime}|\hat{J}^{-}|P\rangle$, must be consistent. The necessity to include zero modes or nonvalence contributions, is often a consequence of the truncation of the Fock space or the specific component of the current employed. The authors in Zhang et al. (2025) (and references therein) investigate how the proper treatment of these terms is essential for the cancellation of spurious frame dependence, a clear signature of the restoration of Poincaré symmetry de Melo et al. (1998, 1999).
@@
Light-front dynamics is inherently a constrained system, and thus precludes direct canonical quantization. This noncanonical structure introduces complications, particularly in fermionic theories, where certain components of the spinor fields are not independent degrees of freedom; they are subject to constraint equations, and these are often referred to as bad components. These bad components necessarily enter into the definition of the electromagnetic operators, subsequently leading to modified, potentially nonstandard currents. Despite the fact that the simplest Ward-Takahashi identity connects the electromagnetic vertex and the propagator, the primary unanswered question involves the integral relationship tying the current to the vertex.
@@
The current investigation significantly advances the initial results by exploring a more generalized scope (or a broader class of systems/parameters). Central to this expanded analysis is a detailed examination of the WT identities as they apply to the fermionic sector of the theory. Crucially, the analysis reaffirms the persistent requirement to incorporate nonvalence contributions—those originating outside the simplest quasiparticle picture, to accurately model the electromagnetic current.
@@
In the present work, with a one-loop current in the Yukawa model, the only dynamical fields in the loop integrals are massive fermions. Consequently, the zero-mode contributions, which are critical for infrared structure and for phenomena like confinement in certain frameworks, stem solely from these particles. Within this framework, the fermion mass $m$ provides a suppression of the fermionic zero-mode effects, thereby yielding a more subdued infrared structure compared to gauge-theoretic scenarios. However, in non-Abelian gauge theories like QCD, the dominant infrared dynamics, including the generation of a confining potential, is primarily driven by the gauge boson zero modes. The gluon, in actuality massless in the Lagrangian, provides unsuppressed contributions in the infrared regime. In the case of the Abelian quantum electrodynamics, the photon zero modes can lead to significant long-range effects, which are absent in purely Yukawa systems.
The present article is organized as follows. In Sec. II, we provide a proof for the Ward-Takahashi identity at one loop in the covariant formulation for the given system. It is followed by the proof in the LF formulation in Sec. III. In Sec. III.1, the LF expression for self-energy is obtained by integrating out the energy component in the loop momentum. The contribution of zero modes to this expression is explained in the process. Further, the LF expression for the electromagnetic current is similarly obtained in Sec. III.2 and the LF Ward-Takahashi identity is established. A brief qualitative discussion about electromagnetic form factor in the context of the issue at hand is given in Sec. III.3. Section IV contains a few concluding remarks. The conventions used throughout the article and a few useful identities are provided in Appendix A, and a part of the algebra from Sec. III.2 is relegated to Appendix B.
@@ II Ward-Takahashi identity in the covariant formulation
Consider a spin-0 bound-state made up of a fermion-antifermion pair, for example a pion, given by a $\gamma^{5}$ type of an interaction vertex. The self-energy diagram for such a state is shown in Fig. 1.
@@
The expression for the self-energy of this bound state is given in the equal-time covariant formulation by
@@
Now consider the electromagnetic current corresponding to this bound state. It is represented by the Feynman triangle diagram shown in Fig. 2.
@@ III Ward-Takahashi identity in the LF formulation
The light-front formulation is usually implemented in the form a Hamiltonian field theory. Thus, all particles in the theory are on shell. We now establish the Ward-Takahashi identity at one-loop for the system in Sec.II i.e. a fermion-antifermion bound-state of spin 0, but in the LF formulation. For this, we consider the 4D loop integral in the covariant theory and integrate out the energy component from the loop integral. This procedure is essentially equivalent to calculating the time-ordered graphs of the theory in old-fashioned Hamiltonian perturbation theory (Bhamre and Misra, 2022). While establishing the Ward-Takahashi identity such, one needs to be mindful of the zero modes that arise when the longitudinal LF momentum vanishes. We proceed to show this below.
@@ III.1 The self-energy
where $k^{-}_{\text{on}}=\frac{k_{\perp}^{2}+m^{2}}{2k^{+}}$ is the on-shell LF energy of a particle with momentum $k^{\mu}=(k^{+},k^{-},k^{\perp})$. Thus, a fermion propagator can be split into an “on-shell“ term and an “off-shell” term as follows
@@
the on-shell $4$-momentum being given by $k^{\mu}_{\text{on}}=(k^{+},k^{-}_{\text{on}},k^{\perp})$. The “off-shell” term is a nonpropagating term that is typically present in the LF formulation. Similarly, the other propagator in the fermion loop is written as
@@
Using the properties of gamma matrices $\{\gamma^{5},\gamma^{\mu}\}=0$ and $(\gamma^{+})^{2}=0$, it is seen that the last term of the above integrand vanishes. Also, the traces in the remaining terms can be calculated using the properties of gamma matrices, giving
@@
$I_{1}^{-}$ in Eq. (16) has poles at $k_{1}^{-}=\frac{f_{1}-i\epsilon}{2k^{+}}$ and $k_{2}^{-}=P^{-}-\frac{f_{2}-i\epsilon}{2(P^{+}-k^{+})}$. Note that there is no $k^{-}$ dependence in the numerator as the $4$-momenta appearing there ($k^{\mu}_{\text{on}}$, $(P-k)^{\mu}_{\text{on}}$) are on-shell momenta. For $k^{+}<0$, both the poles $k_{1}^{-}$ and $k_{2}^{-}$ lie above the real axis while for $k^{+}>P^{+}$ , both lie below it. Hence, closing the contour on the opposite side of the position of the pole, the integral vanishes for each of these two ranges of the longitudinal momentum. For $0<k^{+}<P^{+}$, $k_{1}^{-}$ lies below the real axis whereas $k_{2}^{-}$ lies above. Closing the contour below, the residue theorem gives
@@
However, this is not the case in Eqs.(17) and (18). $I_{2}^{-}$ in Eq. (17) has a pole at $k_{1}^{-}=\frac{f_{1}-i\epsilon}{2k^{+}}$ which does go to infinity as $k^{+}\to 0$. This is a purely divergent zero-mode term that manifests here as an end-point singularity. In order to correctly account for this pole at infinity, we make a change of variable so as to bring the pole to the origin, regularize it, and eventually take the regulator to zero (Bhamre and Misra, 2022). Let $k^{-}=\frac{1}{u}$ in Eq. (17). Then,
@@
The first integral above has poles at $u_{1}=-i\delta$ and $u_{2}=\frac{2k^{+}}{f_{1}-i\epsilon}$. Thus, the integral is nonvanishing only for $k^{+}>0$. Similarly, the second integral contributes only through the region $k^{+}<0$. Applying the residue theorem, and then taking the regulator $\delta$ to zero, we get
@@
Considering $P^{\prime+}>P^{+}$, the integration regions of the log-divergent $k^{+}$ integrals combine to give
@@ III.2 The electromagnetic current
The $P$-dependent and $P^{\prime}$-dependent parts of $q_{\mu}J^{\mu}(P,P^{\prime})$ in Eq. (37) separate out and are contained in the two sets of integrals $I^{+}_{1}$ through $I^{+}_{6}$, and $I^{+}_{7}$ through $I^{+}_{12}$, respectively. Substituting the sum over all $I^{+}_{i}$ in Eq. (37) eventually gives (algebra worked out in Appendix B)
@@
Leaving the zero-mode contribution in Eq. (38) aside, it is worth noting that there are two regions of $k^{+}$ integration-one from 0 to $P^{+}$, and another from $P^{+}$ to $P^{\prime+}$. Considered as a time-ordered graph in LF, the diagram in Fig. 2 can only contribute to the region $0<k^{+}<P^{+}$ due to 3-momentum conservation and the fact that the longitudinal momentum component in LF can only be non-negative. The contribution to the rest of the longitudinal momentum region, i.e., $P^{+}<k^{+}<P^{\prime+}$ is provided by the other time-ordered LF diagram, viz. the “pair production” diagram shown in Fig. 3. Thus, without the pair-production diagram, the Ward-Takahashi identity in LF cannot be established.
@@
The Ward-Takahashi Identity (WTI) is an expression of gauge invariance, and a proper regularization scheme should ideally preserve the WTI exactly. In light-front field theory, the WTI is a statement about the conserved current $J^{\mu}$. Its longitudinal structure (involving $k^{+}$) is typically what determines the existence of zero modes. A transverse ultraviolet cutoff $\Lambda_{\perp}$ (on $k_{\perp}$), provided it is independent of the LF coordinate $x^{+}$ and LF momentum $k^{+}$, is generally compatible with the structure of the light-front Ward-Takahashi identity (LFWTI). The LFWTI, being a fundamental requirement for current conservation, remains exactly satisfied under a simple transverse momentum cutoff $\Lambda_{\perp}$. The existence of zero modes is related to the integration over the longitudinal momentum $k^{+}$. The nonvalence terms are required, independent of the value of $\Lambda_{\perp}$, although their explicit functional forms will involve $\Lambda_{\perp}$.
The zero-mode contributions for the electromagnetic current, $J^{\mu}_{ZM}$, arise from the behavior of the loop integrand as $k^{+}\to 0$,
$J^{\mu}_{\text{zero modes}}\sim\int d^{2}k_{\perp}[\text{Residue at }k^{+}=0].$ (39)
The explicit zero-mode contribution is obtained by performing the $d^{2}k_{\perp}$ integration. Therefore, the numerical magnitude of the zero-mode contribution is sensitive to $\Lambda_{\perp}$. However, regardless of the $\Lambda_{\perp}$ used, the structural role of the zero modes remains the same: it must contribute the necessary term to ensure $q_{\mu}J^{\mu}(P,P^{\prime})=\Sigma(P)-\Sigma(P^{\prime})$ is satisfied. The zero-mode contribution must be included for any finite $\Lambda_{\perp}$ .
In summary, the inclusion of a transverse cutoff $\Lambda_{\perp}$ does not invalidate the necessity of zero modes, but it makes the final calculated value of the current a function of this regulator, as expected.
III.3 Electromagnetic form factor
An important aspect related to the previous discussion is the calculation of electromagnetic form factors for meson bound states within the LF approach. With the calculated matrix elements of the electromagnetic current given in Eq. (5), it is possible to extract the electromagnetic elastic form factors for a bound state of two spin-half particles. This can be done using the covariant Mandelstam formula (Mandelstam, 1955) given by the following expression,
$\langle P^{\prime}|\hat{J}^{\mu}|P\rangle=(P^{\prime\mu}+P^{\mu})F_{M}(q^{2})$ (40)
and the plus or minus component of the electromagnetic current $J^{+},J^{-}$.
However, the calculation of form factors within the LF approach needs to be carried out with care. This is because of certain symmetries like the rotational symmetry being broken, leading to contributions in addition to the valence contributions.
As established in the works by de Melo et al., the LF projection of the electromagnetic current is often frame dependent. By comparing different current components ($J^{+}$ vs. $J^{-}$), de Melo et al. showed that nonvalence pair terms and zero modes are necessary to obtain a frame-independent form factor (de Melo et al., 1999, 2002). Their models successfully describe $F_{\pi}(Q^{2})$ in both spacelike and timelike regions, incorporating vector meson dominance at the quark level (de Melo et al., 2004).
The one-loop contribution to the current in the LF frame is split into the standard (valence) term $J^{\mu}_{\text{val}}$ and the nonstandard terms $J^{\mu}_{\text{non-val}}$. The necessity of $J^{\mu}_{\text{non-val}}$ is to ensure that the LFWTI holds and that the resulting current is fully covariant. Since $J^{\mu}$ must satisfy the WTI for all values of the momentum transfer $q$, the combined contributions of the zero modes and pair-production diagrams, $J^{\mu}_{\text{non-val}}$, must explicitly depend on $q^{2}$.
Specifically, these nonstandard contributions are precisely the terms required to cancel the explicit breaking of the WTI found in $J^{\mu}_{\text{val}}$ alone, thereby restoring covariance. When decomposing the total current $\langle P^{\prime}|\hat{J}^{\mu}|P\rangle$ into its form factor structure, $F(q^{2})$ will contain contributions from all sectors,
$F(q^{2})=F_{\text{val}}(q^{2})+F_{\text{non-val}}(q^{2}).$ (41)
The dependence of $F_{\text{non-val}}$ on $q^{2}$ is fixed by the constraint that the final result for $F(q^{2})$ must match the covariant calculation, which, by definition, has a nontrivial dependence on $q^{2}$. Thus, the zero-mode and pair-production terms are necessarily $q^{2}$ dependent to ensure that the final result gives the correct form factor (see Refs.(de Melo et al., 1999, 2002; Bakker et al., 2001)). It is worth noting that the zero-mode problem in the light-front approach is not restricted only to spin-half particles. For example, a good recent discussion of zero modes for a Yukawa scalar theory is given in Ref.(Zhang et al., 2025).
Within the light-front quark model, the complete electromagnetic current matrix elements are derived by combining valence and nonvalence contributions. The inclusion of the latter, which account for zero-mode effects, allows for the full restoration of covariant electromagnetic form factors.
@@ IV Conclusion
The preceding analysis delivers a rigorous, one-loop proof of the Ward-Takahashi identity for a spinless, fermionic bound state, comparing and contrasting its realization within both the covariant and light-front formulations of quantum field theory.
@@
By correctly incorporating these end-point singularities, the contributions typically associated with zero modes - the dynamic complexity unique to the light-front - are naturally and elegantly taken into account in the proof. This demonstrates that the WT identity, and thus gauge invariance, is preserved in the LF theory without the need for ad hoc zero-mode operators, provided the loop integrals are handled with scrupulous care. Along with the “regular” diagram viz. the triangle diagram for the electromagnetic current, the “pair production” diagram is also necessary to establish the proof of the identity. This is consistent with the fact that the proof is achieved in the Hamiltonian time-ordered perturbation theory on the light front.
This work not only validates the light-front hamiltonian time-ordered perturbation theory as a consistent framework for handling gauge theories but also provides a clear, instructive procedure for translating between covariant and LF results. The explicit connection made between end-point singularities in loop integrals and the required physical contributions (like pair production diagrams) significantly enhances our understanding of how fundamental symmetries, like gauge invariance, are preserved in nontrivial field-theoretic contexts within the light-front framework. This rigorous verification strengthens the theoretical foundation for future applications of light-front quantization to more realistic and complex bound-state problems in particle and nuclear physics.
Acknowledgments
D. B. acknowledges the support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil, Process No. 152348/2024-7. J. P. B. C. M. acknowledges the support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), Process No. 2023/09539-1, Instituto Nacional de Ciência e Tecnologia-Nuclear Physics and Applications (INCT-FNA, MCTI), Brazil, Process No. 464898/2014-5, and FAPESP, Brazil, thematic Project, No. 2024/17816-8.